What will be the freezing point of an aqueous solution containing 80.0 g of NaCl in 3.4 L of H2O?
1
Expert's answer
2020-10-26T13:55:15-0400

Q139289


What will be the freezing point of an aqueous solution containing 80.0 g of NaCl in 3.4 L of

H2O?


Solution :

Depression in freezing point: The Freezing point of the solvent is lowered on the addition of a non-volatile compound.


It is a colligative property that is it depends on the number of the solute particles present and not on their identity.


The formula for depression in freezing point is


ΔTF = - i * m * KF, ---------------- Equation 1


where, ΔTf is the depression in freezing point

i = van’t hoff factor of the solute.

m = molality of the solution.

Kf = Freezing point depression constant.


We will consider that NaCl dissociates completely in water to give Na+ (aq) and Cl- (aq).

Since each NaCl can given two ions so van’t hoff factor, i = 2.


For water, KF = 1.86 0C/m


Step 1 : To find the moles of NaCl


Convert 80.0g of NaCl to moles using molar mass of NaCl.


Molar mass of NaCl = 1 * atomic mass of Na + 1 * atomic mass of Cl

= 1 * 22.99g/mol + 1* 35.453g/mol

= 58.443g/mol



moles of NaCl = 80.0g of NaCl * 1 mole NaCl/58.443g of NaCl = 1.369 mol NaCl


Step 2: To find the mass of given 3.4L water.


We are given volume of water = 3.4L

Let us assume that the density of water is 1.00 kg/L


Substitute volume and density of water in formula of density and find the mass of water.


Density = mass/ Volume ;

So, mass = Density * volume = 1.00kg/L * 3.4L = 3.4 kg



Step 3: To find the molality of the solution.


Moles of NaCl = 1.369moles and mass of solvent (water ) = 3.4kg



molality = moles of solute/mass of solvent in ‘kg’

molality = 1.369mol/3.4kg = 0.4026m



Step 4: To find the depression in the freezing point


Substitute i = 2, m = 0.4026m and KF = 1.86 0C/m in the Equation 1, we have



ΔTF = - i * m * KF = - 2 * 0.4026m * 1.86 0C/m

ΔTF = - 1.498 0C



Step 5: To find freezing point of water after addition of NaCl .

We know Freezing point of pure water = 0.000C


ΔTF = Tsolution - Twater


-1.498 0C = Tsolution - 0.000C


Tsolution = -1.4980C


In the question the quantity with the least number of significant figure is 3.4L. It has got 2 significant figures. So we must round up our final answer to 2 significant figures.


Hence the freezing point of the aqueous solution will be -1.50C.